OpenStudy (anonymous):

when ordering a complete meal, you have a choice of one main course, one beverage, and one dessert. How many different combinations of complete meals do you have to choose from? Main Course Beverage Desert ________________________________________________________________ Fish Milk Pie Chicken Juice Cookies Salad Bar Cake

OpenStudy (anonymous):

multiply the number of choices together

OpenStudy (phoenixfire):

Main course has 3 options Beverage has 2 options Desert has 3 options Complete combinations would just be 3*2*3.

OpenStudy (anonymous):

\[3\times 2\times 3\]

OpenStudy (anonymous):

23

OpenStudy (phoenixfire):

@smartboy23 Not quite. 3*2=6... 6*3=18

OpenStudy (anonymous):

@PhoenixFire

OpenStudy (phoenixfire):

Yes?